The sum of the ages of Mr. lee and his son is 53 now. 4 years ago, the age of Mr. Lee was four times that of his son. Find their ages now.
Answer by ikleyn(52803)

.
(1)   L + S = 53      ("The sum of the ages of Mr. lee and his son is 53 now")
(2)   L-4 = 4*(S-4)   ("4 years ago, the age of Mr. Lee was four times that of his son")


From equation (1), express  L = 53-S  and substitute it into equation (2). You will get


(53-S) - 4 = 4*(S-4)

49 - S = 4S - 16

49 + 16 = 4S + S

65 = 5S  ====>  S = 65/5 = 13.


Answer.  The sun is 13 years old.  The father (= Mr. Lee) is 53-13 = 40 years old.

Solved.